Given the vectors A=4î + 8j and B= 5î + 10j, demostrate that the vectors are parallels.
Answer #1

When two vectors are parallel, angle between the vectors is zero.

old{A}=4hat{i}+8hat{j}   old{B}=5hat{i}+10hat{j}

old{A}cdot old{B}=ABcos heta

A B AB cos θ

{old{A}cdot old{B}}=(4hat{i}+8hat{j})cdot(5hat{i}+10hat{j})=4*5+8*10=100

A=sqrt{4^2+8^2}=sqrt{80}   B=sqrt{5^2+10^2}=sqrt{125}

cos 1 A AB

C125 cos11) COS

Angle between the two vectors is 0degree , so the given vectors are parallel.